A train 300m. Long passed a man walking along the line in the same direction at the rate of 3 km./hr in 33 second . The speed of train?

Grade:12

1 Answers

Agrim
6 years ago
D=300m S of man =3 km/h=3*5/18 m/s =5/6 m/sT=33 sD covered by man in 33 s=s*t =5/6*33 =165/6 mD covered by train per 1 m of man`s path=300 mD covered by train to pass whole path of man=165/6*300 =8250 m =8.25 kmS of train=D/T =8250/33 =250 m/s =250*18/5 =50*18 =900 km/h